# STORE: Streamlining Semantic Tokenization and Generative Recommendation with A Single LLM

### v1

@@ 1. Introduction
Recommender systems are widely deployed in online applications, such as e-commerce websites, advertising networks, streaming services, and social media, to deliver personalized recommendations tailored to users’ interests. However, traditional recommendation models often rely on unique item identifiers (IDs) to represent items (Yuan et al., 2023; Liu et al., 2024b), which presents several key limitations. First, ID-based item representation can suffer from overfitting due to the typically sparse and imbalanced nature of the training data. Second, it fails to fully leverage item content information, which is crucial for improving recommendations for long-tail and cold-start items. Third, ID embeddings are learned and updated independently, making it difficult to capture the semantic relationships among items. Finally, given the vast number of items, the item embedding table often consumes a significant amount of memory.
To address these limitations, semantic tokenization has recently emerged as a promising solution and has gained rapid traction in the community (Singh et al., 2024; Rajput et al., 2024; Liu et al., 2024a; Zhu et al., 2024). As illustrated in Figure 1, instead of representing each item with a unique ID embedding, semantic tokenization encodes each item’s semantic representation into a compact sequence of discrete tokens. These tokens, also known as semantic identifiers or codes [^1] In this paper, the terms semantic identifiers, discrete tokens, and codes may be used interchangeably., preserve the item’s semantics into the token space and ensure that semantically similar items are represented by similar tokens. For example, consider a scenario where each item is represented by a sequence of 4 consecutive tokens, each capable of assuming one of 256 possible values. This theoretically creates a representation space of approximately $256^{4}\approx 4$ billion combinations, which is sufficient or easily expandable for industrial-scale systems. Additionally, tokens and their embeddings can be shared across items, allowing the similarity between two items to be roughly estimated by the Hamming distance between their token sequences. This information can be obtained without the need for training on subsequent recommendation tasks and is independent of token embeddings. As a result, semantic tokenization effectively addresses the aforementioned challenges and shows promise in semantic-enhanced modeling for recommendations (Rajput et al., 2024; Singh et al., 2024).
@@ 2. Preliminaries
Traditional recommender systems use unique item identifiers mapped to learnable vectors (embeddings) to represent items during training, leading to the challenges mentioned earlier. Recent studies (Rajput et al., 2024; Zheng et al., 2024; Liu et al., 2024a) have shown that semantic tokenization is becoming an increasingly popular approach for embedding item content into identifiers. [^2] Recent works (Wang et al., 2024b, a; Qu et al., 2024) have extracted item embeddings from pretrained recommenders using collaborative knowledge. However, these collaborative embeddings are unstable and frequently change in real-world scenarios, unlike the stable nature of item content knowledge. Therefore, this paper focuses on generating semantic identifiers merely based on item content features. In this section, we focus on two key aspects: i) generating semantic identifiers using standard approaches, and ii) utilizing these semantic identifiers in downstream recommendation models.
@@ 5.1. LLMs for Recommendation
Research in this paradigm typically involves tasks designed to model diverse user behaviors and aims to develop a fundamental recommendation model. For instance, PTUM (Wu et al., 2020b) employs two pre-training tasks: masked behavior prediction and next K behavior prediction. Similarly, Cui et al. (Cui et al., 2022) introduce M6, which utilizes an auto-regressive generation task and a text-infilling objective. Additionally, Geng et al. (Geng et al., 2022) propose P5, a model that integrates multiple recommendation tasks within a unified framework to pre-train a foundational recommendation model.
@@
Instead of pre-training an LLM, some studies aim to directly integrate LLMs into the recommendation pipeline without parameter updates, typically through feature augmentation. For instance, Xi et al. (Xi et al., 2023) propose utilizing LLMs to infer user preferences and factual knowledge about items. Similarly, Wang et al. (Wang and Lim, 2023) employ LLMs to model user preferences.
@@ 5.2. Generative Recommendation
To incorporate item content knowledge, TIGER (Rajput et al., 2024) introduced semantic identifiers that can be shared across different items, replacing the unique identifier. This approach has been further developed by other semantic tokenization efforts (Zheng et al., 2024; Jin et al., 2023; Liu et al., 2024a). Moreover, the effectiveness of integrating collaborative features learned from simple recommenders into identifiers has been demonstrated (Qu et al., 2024; Wang et al., 2024a). However, this method heavily relies on rich interactions and tends to be unstable or change frequently in real-world scenarios. The learned tokens are used in generative recommenders, shifting the training focus from next-item prediction to next-code prediction. This shift narrows the search space for each position, thereby enhancing the inference performance. Given the limitations discussed in the Preliminaries Section, this paper aims to reevaluate and refine the standard semantic tokenization pipeline.

### v2

@@ 1. Introduction
Recommender systems are widely deployed in online applications, such as e-commerce websites, advertising networks, streaming services, and social media, to deliver personalized recommendations tailored to users’ interests. However, traditional recommendation models often rely on unique item identifiers (IDs) to represent items (Yuan et al., 2023; Liu et al., 2024d), which presents several key limitations. First, ID-based item representation can suffer from overfitting due to the typically sparse and imbalanced nature of the training data. Second, it fails to fully leverage item content information, which is crucial for improving recommendations for long-tail and cold-start items. Third, ID embeddings are learned and updated independently, making it difficult to capture the semantic relationships among items. Finally, given the vast number of items, the item embedding table often consumes a significant amount of memory.
To address these limitations, semantic tokenization has recently emerged as a promising solution and has gained rapid traction in the community (Singh et al., 2024; Rajput et al., 2024; Liu et al., 2024b; Zhu et al., 2024). As illustrated in Figure 1, instead of representing each item with a unique ID embedding, semantic tokenization encodes each item’s semantic representation into a compact sequence of discrete tokens. These tokens, also known as semantic identifiers or codes [^1] In this paper, the terms semantic identifiers, discrete tokens, and codes may be used interchangeably., preserve the item’s semantics into the token space and ensure that semantically similar items are represented by similar tokens. For example, consider a scenario where each item is represented by a sequence of 4 consecutive tokens, each capable of assuming one of 256 possible values. This theoretically creates a representation space of approximately $256^{4}\approx 4$ billion combinations, which is sufficient or easily expandable for industrial-scale systems. Additionally, tokens and their embeddings can be shared across items, allowing the similarity between two items to be roughly estimated by the Hamming distance between their token sequences. This information can be obtained without the need for training on subsequent recommendation tasks and is independent of token embeddings. As a result, semantic tokenization effectively addresses the aforementioned challenges and shows promise in semantic-enhanced modeling for recommendations (Rajput et al., 2024; Singh et al., 2024).
@@ 2. Preliminaries
Traditional recommender systems use unique item identifiers mapped to learnable vectors (embeddings) to represent items during training, leading to the challenges mentioned earlier. Recent studies (Rajput et al., 2024; Zheng et al., 2024; Liu et al., 2024b) have shown that semantic tokenization is becoming an increasingly popular approach for embedding item content into identifiers. [^2] Recent works (Wang et al., 2024b, a; Qu et al., 2024) have extracted item embeddings from pretrained recommenders using collaborative knowledge. However, these collaborative embeddings are unstable and frequently change in real-world scenarios, unlike the stable nature of item content knowledge. Therefore, this paper focuses on generating semantic identifiers merely based on item content features. In this section, we focus on two key aspects: i) generating semantic identifiers using standard approaches, and ii) utilizing these semantic identifiers in downstream recommendation models.
@@ 5.1. LLMs for Recommendation
(Liu et al., 2022; Wu et al., 2020b; Cui et al., 2022; Geng et al., 2022; Liu et al., 2024c) Research in this paradigm typically involves tasks designed to model diverse user behaviors and aims to develop a fundamental recommendation model. For instance, PTUM (Wu et al., 2020b) employs two pre-training tasks: masked behavior prediction and next K behavior prediction. Similarly, Cui et al. (Cui et al., 2022) introduce M6, which utilizes an auto-regressive generation task and a text-infilling objective. Additionally, Geng et al. (Geng et al., 2022) propose P5, a model that integrates multiple recommendation tasks within a unified framework to pre-train a foundational recommendation model.
@@
(Xi et al., 2023; Wang and Lim, 2023; Liu et al., 2024a) Instead of pre-training an LLM, some studies aim to directly integrate LLMs into the recommendation pipeline without parameter updates, typically through feature augmentation. For instance, Xi et al. (Xi et al., 2023) propose utilizing LLMs to infer user preferences and factual knowledge about items. Similarly, Wang et al. (Wang and Lim, 2023) employ LLMs to model user preferences.
@@ 5.2. Generative Recommendation
To incorporate item content knowledge, TIGER (Rajput et al., 2024) introduced semantic identifiers that can be shared across different items, replacing the unique identifier. This approach has been further developed by other semantic tokenization efforts (Zheng et al., 2024; Jin et al., 2023; Liu et al., 2024b). Moreover, the effectiveness of integrating collaborative features learned from simple recommenders into identifiers has been demonstrated (Qu et al., 2024; Wang et al., 2024a). However, this method heavily relies on rich interactions and tends to be unstable or change frequently in real-world scenarios. The learned tokens are used in generative recommenders, shifting the training focus from next-item prediction to next-code prediction. This shift narrows the search space for each position, thereby enhancing the inference performance. Given the limitations discussed in the Preliminaries Section, this paper aims to reevaluate and refine the standard semantic tokenization pipeline.